OpenStudy (anonymous):

i try to solve this in fractions but i had a hard time to know this i tried solving it but my answer's are wrong when i check it help me.......!!! x-3y=7 and 4x+y=6 in substitution way.

OpenStudy (anonymous):

solve for x in the first equation: x-3y = 7 making x = 7 + 3y then, substitute the value of x to the 2nd equation: 4(7+3y) + y = 6. Now try to solve it.